Total Cost Concept of Product Pricing

Voice Com, Inc., produces and sells cellular phones. The costs of producing and selling 4,000 units of cellular phones are as follows:

Variable costs: Fixed costs:
    Direct materials $ 74 per unit     Factory overhead $148,000
    Direct labor 34     Selling and admin. exp. 52,000
    Factory overhead 22
    Selling and admin. exp. 18
         Total $148 per unit

Voice Com desires a profit equal to a 16% rate of return on invested assets of $386,100.

Assume that Voice Com, Inc., uses the total cost concept of applying the cost-plus approach to product pricing.

a. Determine the total costs and the total cost amount per unit for the production and sale of 4,000 units of cellular phones. Round the cost per unit to two decimal places.

Total cost $
Cost amount per unit $

b. Determine the total cost markup percentage (rounded to two decimal places) for cellular phones.
%

c. Determine the selling price of cellular phones. Round to the nearest cent.
$per phone
Answer #1

Total Variable Cost = $148 * 4000 units = $592000

Total Fixed Cost = $148000 + $52000 = $ 200000

Required rate of return = $386100 * 16% = $61776

a.1) Total Cost = $592000 + $200000 = $792000

2) Cost per unit = $792000/4000 = $198

b. Total Cost Markup % = $61776/ $792000 = 7.80 %

c. Selling Price = (Total Cost + Profit) / Units

(792000 + 61776) / 4000 = $213.44